OpenStudy (jennithemeani):

A poll was conducted in states A, B, and C to determine which candidate voters would most likely vote for. The results are displayed in the table below: State A State B State C Total Liberal 22 23 13 58 Conservative 18 12 12 42 Total 40 35 25 100 What is P(liberal | State A)?

OpenStudy (p0sitr0n):

use Bayes rule: P(A|B)=P(A,B)/P(B). Here, P(L|A)=22/80=.275

OpenStudy (jennithemeani):

Where'd you get the 80 from?

OpenStudy (jennithemeani):

nvm, I think I got it